The most recent financial statements for Moose Tours, Inc., appear below. Sales for 2016 are projected to grow by 20 percent. Interest expense will remain constant; the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales MOOSE TOURS, INC 2015 Income Statement Sales Costs Other expenses $755,000 590,000 11,000 Earnings before interest and taxes Interest expense $154,000 12,000 Taxable income Taxes (40%) $142,000 56,800 Net income $ 85,200 Dividends Addition to retained earnings $34,080 51,120 MOOSE TOURS, INC Balance Sheet as of December 31, 2015 Assets Liabilities and Owners Equity Current assets Current liabilities Cash Accounts receivable Inventory $ 21,440 Accounts payable 33,760 Notes payable $ 55,600 14,800 $70,400 $138,000 720 Total Tota $125,920 Long-term debt Fixed assets Owners equity Net plant and equipment 270,000 Common stock and paid-in surplus $124,000 63,520 $187,520 $395,920 Retained earnings Tota Total assets $395,920 Total liabilities and owners equityWhat is the EFN if the firm was operating at only 80 percent of capacity in 2015? Assume that fixed assets are sold so that the company has a 100 percent asset utilization. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g, 32 intermediate whole number, e.g., 32.) EFN
Answer #1

EFN = Increase in assets – Increase in liabilities – Increase in retained earnings

Step 1:

Increase in retained earnings

Sales = 755000*120%= 906000

Less: costs 590000*120%=708000

Other expenses = 11000*120%= 13200

EBIT = 184800

Less: Interest 12000

Taxable income = 172800

Tax = 40% = 69120

Net Income = 103680

Dividend = (34080* 103680/85200)= 41472

Addition to retained earnings = 62208

Step 2: Increase in assets = (Current + Fixed assets)*20%

= (125920+270000)*20%

=79184

Increase in spontaneous liabilities = Accounts payable *20%

=55600*20%

=11120

Step 3:

EFN = Increase in assets – Increase in liabilities – Increase in retained earnings

EFN = 79184-11120- 62208

=5856